OpenStudy (anonymous):

How would I graph the function \[f(x)=\sqrt{x-1}-1\] ? The directions say to start with the leftmost point and graph three additional ones. Help?

jimthompson5910 (jim_thompson5910):

x = 1 is the smallest whole number you can plug in, so let's start there

jimthompson5910 (jim_thompson5910):

\[f(x)=\sqrt{x-1}-1\] \[f(1)=\sqrt{1-1}-1\] \[f(1)=\sqrt{0}-1\] \[f(1)=0-1\] \[f(1)=-1\] So the point (1, -1) is on the graph

jimthompson5910 (jim_thompson5910):

Let's now plug in x = 2 \[f(x)=\sqrt{x-1}-1\] \[f(2)=\sqrt{2-1}-1\] \[f(2)=\sqrt{1}-1\] \[f(2)=1-1\] \[f(2)=0\] So the point (2, 0) is on the graph

jimthompson5910 (jim_thompson5910):

I'll let you do the other two points